OpenStudy (anonymous):

A triangle has coordinates A (1, 5), B (-2, 1) and C (0, -4). What are the new coordinates if the triangle is dilated with a scale factor of 1/5?

OpenStudy (anonymous):

A=(1/5,1) B=(-2/5,1/5) C=(0,-4/5)

OpenStudy (anonymous):

Simply multiply scale factor by each of the coordinates

OpenStudy (anonymous):

ok thanx